OpenStudy (anonymous):

x^2+7x=0 Solve please?(:

OpenStudy (anonymous):

HInt: Take x outside..:)

OpenStudy (anonymous):

Whaa??

OpenStudy (anonymous):

x(x+7)=0 Now solve

OpenStudy (anonymous):

x^2-7x-10=0 splitting the middle term we get, x^2-5x-2x-10=0 x(x-5)-2(x-5)=0 (x-2) (x-5)=0

OpenStudy (anonymous):

Thats the answer JJ;)

OpenStudy (anonymous):

You Understand???

OpenStudy (anonymous):

I am.. SO confused..

OpenStudy (anonymous):

x=7

OpenStudy (anonymous):

my answers are x=0 or 7 x=7 or -7 and x-0 or -7

OpenStudy (anonymous):

its x=0 or 7

OpenStudy (anonymous):

Solve x(x+7) =0

OpenStudy (anonymous):

x=7 or -7 is your answer for sure of it.

OpenStudy (anonymous):

Nope. No it's not. Y'all are both wrong. It was x=0 or -7 @Champs is it even possible to solve for that?

OpenStudy (anonymous):

Um ok well idk how those are the possible outcomes theres no way any of those are right.

OpenStudy (anonymous):

Well.. I don't know.

OpenStudy (anonymous):

\[x^{2}+7\] can be written as x( x+7) =0 right? For this equation to be zero either x=0 or x+7 =0. So you will get x=0 or X= -7. Clear? @YouCanCallMeJJ

OpenStudy (anonymous):

Clear as a button :D

OpenStudy (anonymous):

Lol..:D